OpenStudy (anonymous):

Please Check! Line integral: \[\int\limits_{C}^{}x ^{2}dx+y ^{2}dy\] of the line segment from (0,2) to (4,3)

OpenStudy (anonymous):

Is it: \[\int\limits_{0}^{4}16t ^{2}(4)+(2+t)^{2}(1)dt\]

OpenStudy (anonymous):

The parameterized curve is r(t)=(4t, 2+t) right?

OpenStudy (anonymous):

@phi @amistre64 @Zarkon @TuringTest

OpenStudy (amistre64):

id have a little reading to do to catch up on this :)

OpenStudy (anonymous):

I think I have it down correctly, I'm just not sure if the limit is from 0 to 4

OpenStudy (amistre64):

we need to parametrize the line from 0,2 to 4,3 right?

OpenStudy (anonymous):

Yes

OpenStudy (amistre64):

4,3 -0-2 ----- <4,1> is the direction vector, and we can apply it to either point given so one parametric could be: x = 0 + 4t y = 2 + t

OpenStudy (anonymous):

That's what I got! I'm not sure about the limits of integration though

OpenStudy (abb0t):

sketch out the graph.

OpenStudy (amistre64):

the dx and dy parts in your post got me confuddled. im used to seeing a ds in it

OpenStudy (amistre64):

\[\int_CPdx+Qdy=\int_CPdx+\int_CQdy\]

OpenStudy (amistre64):

x = 0 + 4t dx = 4 dt y = 2 + t dy = dt

OpenStudy (anonymous):

I have that down @amistre64 I'm just really not sure about the limits. I just used the x coordinates.

OpenStudy (amistre64):

what are the limits of t for the given line?

OpenStudy (anonymous):

1 and 4?

OpenStudy (amistre64):

when t=0 we are at (0,2) when t=? we are at (4,3)

OpenStudy (anonymous):

I'm not sure?

OpenStudy (amistre64):

0 = 0 + 4t 2 = 2 + t t=0 4 = 0 + 4t 3 = 2 + t t = 1 so the limit of our t is 0 to 1

OpenStudy (anonymous):

Wait how did you do that?

OpenStudy (amistre64):

we are changing x and y into functions of t correct?

OpenStudy (anonymous):

Yes

OpenStudy (amistre64):

then "with respect to t" the line is from t=0 to t=1 since (0,2) is t=0; and (4,3) is t=1

OpenStudy (anonymous):

I see!

OpenStudy (amistre64):

we are simply defining all the terms with respect to t; x,y and movement

OpenStudy (anonymous):

I have another question that asks for the line integral that consists of the arc of the circle x^2+y^2=4 from (2,0 to (0,2) I know how to find the line integral, but the limits are what I'm not sure about again

OpenStudy (anonymous):

The parameterized curve is (2cost,2sint)

OpenStudy (amistre64):

is it asking you to define it by the line that goes from 2,0 to 0,2? or from the arc that goes from 2,0 to 0,2 ?

OpenStudy (anonymous):

It says the arc of the circle from (2,0) to (0,2)

OpenStudy (amistre64):

x = r cos(t) y = r sin(t) and from the equation r = 2 x = 2 cos(t) dx = -2sin(t) y = 2 sin(t) dy = 2 cos(t) what is the value of t for the point x=2, y=0? what is the value of t for the point x=0, y=2?

OpenStudy (anonymous):

cos(t)=1 and sin(t)=1?

OpenStudy (amistre64):

2 = 2 cos(t) 0 = 2 sin(t) cos(t) = 1 and sin(t) = 0 when t = 0 0 = 2 cos(t) 2 = 2 sin(t) cos(t) = 0 and sin(t) = 1 when t = pi/2

OpenStudy (anonymous):

So it's from 0 to pi/2 right?

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

Thank you so much!!

OpenStudy (amistre64):

youre welcome, and good luck :)